# : Generalizing Large Language Models for E-commerce

### v1

@@ Abstract
With tremendous efforts on developing effective e-commerce models, conventional e-commerce models show limited success in generalist e-commerce modeling, and suffer from unsatisfactory performance on new users and new products β a typical out-of-domain generalization challenge. Meanwhile, large language models (LLMs) demonstrate outstanding performance in generalist modeling and out-of-domain generalizability in many fields. Toward fully unleashing their power for e-commerce, in this paper, we construct , the first open-sourced, large-scale, and high-quality benchmark instruction dataset for e-commerce. Leveraging , we develop , a series of e-commerce LLMs, by instruction-tuning general-purpose LLMs. Our comprehensive experiments and evaluation demonstrate that models substantially outperform baseline models, including the most advanced GPT-4, and the state-of-the-art task-specific models in in-domain evaluation. Moreover, exhibits excellent generalizability to out-of-domain settings, including unseen products and unseen instructions, highlighting its superiority as a generalist e-commerce model. Both the dataset and the models show great potential in empowering versatile and effective LLMs for e-commerce. and models are publicly accessible through https://ninglab.github.io/eCeLLM.
@@ 1 Introduction
The Internet's evolution and the rise of the digital economy have made e-commerce an integral part of daily life, drawing considerable attention from researchers. In recent years, tremendous research efforts have been dedicated to developing effective e-commerce modelsΒ (Yang etΒ al., 2022; Geng etΒ al., 2022). Though promising, conventional e-commerce models generally suffer from two issues. (1) Limited success in generalist e-commerce modelingΒ (Zhang etΒ al., 2023b): Conventional e-commerce models are typically task-specific (e.g., gSASRec for sequential recommendationΒ (Petrov & Macdonald, 2023), SUOpenTag for attribute value extractionΒ (Xu etΒ al., 2019)). However, contemporary e-commerce platforms, such as Amazon and eBay, are highly complex and consistently expanding, with many interdependent tasks. In this context, generalist modeling is particularly suitable and desired on these platforms due to its cost-effectiveness and extensibility, while the task-specific modeling scheme struggles with scalability. (2) Unsatisfactory performance on new users and new productsΒ (Hou etΒ al., 2023): Cold start (i.e., performing e-commerce tasks on new users or new products)Β (Lika etΒ al., 2014), the typical out-of-domain (OOD) generalization challenge in e-commerce, has been a long-standing and difficult problemΒ (Ding etΒ al., 2022; Yang etΒ al., 2022). Existing task-specific e-commerce models are typically tailored to existing users and products and lack the ability to effectively extrapolate to new users and new productsΒ (Lika etΒ al., 2014). However, new users and new products are very common and highly wanted in the dynamic landscape of e-commerce.
Recently, large language models (LLMs), such as GPT-4Β (OpenAI, 2023), Claude 2Β (Anthropic, 2023), GeminiΒ (Team etΒ al., 2023), and Llama 2Β (Touvron etΒ al., 2023b), have demonstrated exceptional performance in natural language processingΒ (Zhao etΒ al., 2023), information retrievalΒ (Spatharioti etΒ al., 2023), and many other fieldsΒ (Frieder etΒ al., 2023; Zeng etΒ al., 2023). However, their power for e-commerce is not fully unleashed. While limited effortsΒ (Yang etΒ al., 2023; Zhang etΒ al., 2023a; Li etΒ al., 2023b) have been dedicated to leveraging LLMs for e-commerce, most of them focus on studying the utility of pre-trained LLMs in single or homogenous tasks (e.g., e-commerce authoring). Some recent efforts employ instruction tuning to adapt LLMs for e-commerce applicationsΒ (Shi etΒ al., 2023; Li etΒ al., 2023b; Geng etΒ al., 2022). However, they fall short in the covered e-commerce tasks and instruction data. We aim to bridge the gap and develop e-commerce foundation models with real-world utilities for a large variety of e-commerce applications.
@@
Leveraging , we develop a series of e-Commerce LLMs, denoted as (pronounce: e-sell em, /I-βsel@m/) models, by instruction-tuning 6 general-purpose LLMs, such as Llama 2Β (Touvron etΒ al., 2023b) and MistralΒ (Jiang etΒ al., 2023). The models are extensively evaluated across various settings, including IND and OOD data, unseen instructions, and different training sample sizes of . Overall, our experimental results demonstrate the following findings:
@@
FigureΒ 1 depicts the overall scheme of . To the best of our knowledge, this study is the first comprehensive and systematic study of instruction-tuning LLMs for e-commerce applications, open-sources the first large-scale, high-quality benchmark dataset (), and develops the state-of-the-art generalist LLMs ( series) for e-commerce. and models are publicly accessible through https://ninglab.github.io/eCeLLM.
@@ 2 Related Work
Instruction Tuning with LLMsΒ Β Instruction tuning enables the transfer of general knowledge captured in LLMs to specific application domains, facilitating the generalizability of LLMs. FLANΒ (Wei etΒ al., 2021) and T0Β (Sanh etΒ al., 2021) are the early work to investigate instruction tuning by evaluating the zero-shot performance of fine-tuned LLMs on numerous datasets. Both of the models show encouraging performance over the original GPT-3Β (Brown etΒ al., 2020), indicating the effectiveness of instruction tuning. The importance of datasets, model scale, and instructions for instruction tuning are explored in a later work FLAN-v2Β (Chung etΒ al., 2022). Considering the significant impact of datasets for instruction tuning, several studies collect sizable benchmarks consisting of numerous tasks, such as Super-NaturalInstructionsΒ (Wang etΒ al., 2022b), Self-instructΒ (Wang etΒ al., 2022a) and Flan CollectionΒ (Longpre etΒ al., 2023), for fine-tuning general-purpose LLMs on NLP tasks. Meanwhile, instruction-tuned LLMs are employed in specific domains. For example, PlatypusΒ (Lee etΒ al., 2023a) is fine-tuned for reasoning in STEM, and MammothΒ (Yue etΒ al., 2023) is for general math problem-solving. Our incorporates instruction tuning with LLMs for e-commerce.
LLMs for E-commerceΒ Β LLMs are emerging in the general e-commerce realm. RecMindΒ (Wang etΒ al., 2023) utilizes LLMs as an autonomous recommender agent by enhancing their capability to comprehend user behaviors. LLaMA-EΒ (Shi etΒ al., 2023) is fine-tuned on LLaMAΒ (Touvron etΒ al., 2023a) for various e-commerce authoring tasks, such as ads generation and general product QA. EcomGPTΒ (Li etΒ al., 2023b) is a pioneer instruction-tuned LLM on its dataset EcomInstruct. However, most of its tasks are human-constructed by repurposing data from other tasks (e.g., from the ``query product matching" task, a task is constructed to generate user queries based on the matched products), and thus, could have limited utility in real applications. A recent studyΒ (Geng etΒ al., 2022) develops a fine-tuned LLM, P5, with personalized prompts to conduct a few specific recommendation tasks (e.g., sequential recommendation). These studies demonstrate the utility of LLMs and instruction tuning in e-commerce applications. To the best of our knowledge, is the most comprehensive instruction-tuned e-commerce LLM.
Comparison among Instruction-tuned LLMs for E-commerceΒ Β TableΒ 1 summarizes the difference of our from LLaMA-E and EcomGPT, the two existing fine-tuned LLMs capable of performing multiple e-commerce tasks (P5 only performs recommendation). Fundamentally different from LLaMA-E and EcomGPT, focuses on comprehensive real-world e-commerce tasks and data, includes very extensive evaluation and benchmarking, and open-sources instruction data and models.
@@
| Β (Shi etΒ al., 2023) | Β (Li etΒ al., 2023b) | (ours) | |
Instruction tuning |
β | β | β |
@@
Real-world tasks |
β | β | β |
Β Β Β Β Β # training tasks |
5 | 122 | 10 |
Β Β Β Β Β # In-domain tests |
5 | 0 | 10 |
Β Β Β Β Β # out-of-domain tests |
0 | 12 | 6 |
# general-purpose |
5 | 3 | 5 |
LLMs evaluated |
# task-specific SoTA models evaluated |
0 | 0 | 11 |
@@
EcomGPT has 122 training tasks, most of which are manipulated from data of different other tasks. The training data is not publicly available. EcomGPT releases 12 test tasks (8 in Chinese) for only out-of-domain evaluation.
@@ 3 Dataset
We introduce , an instruction dataset for adapting LLMs to e-commerce tasks. features 3 key design principles. (1) Broad coverage: includes 10 diverse tasks of 4 categories. Comprehensive and wide-ranging related tasks are critical in enabling versatile LLMs for a specific domain as shown in the literatureΒ (Lee etΒ al., 2023b). (2) Realistic tasks: focuses on real-world e-commerce tasks with real-world data, not human-manipulated tasks with synthetic data. This ensures e-commerce LLMs tuned on a potentially high utility in real-world applications. (3) High quality: considers only real-world data and undergoes rigorous scrutiny to ensure its accuracy and high quality. High data quality plays a pivotal role in building effective LLMsΒ (Hoffmann etΒ al., 2022; Gadre etΒ al., 2023).
@@ 3.1 E-commerce Tasks
Particularly, includes 3 tasks for product understanding: (1) attribute value extraction ()Β (Xu etΒ al., 2019; Wang etΒ al., 2020; Yang etΒ al., 2022), (2) product matching ()Β (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) and (3) product relation prediction ()Β (Ahmed etΒ al., 2021; Xu etΒ al., 2020). For user understanding, includes (4) sentiment analysis ()Β (Wankhade etΒ al., 2022) and (5) sequential recommendation ()Β (Petrov & Macdonald, 2023; Li etΒ al., 2023). also covers 3 query product matching tasksΒ (Reddy etΒ al., 2022): (6) multi-class product classification (), (7) product substitute identification (), and (8) query-product ranking (). For product question answering, contains the tasks of (9) answerability prediction ()Β (Gupta etΒ al., 2019) and (10) answer generation ()Β (Deng etΒ al., 2023). More details are available in AppendixΒ A. TableΒ A1 summarizes the tasks and their data sources. Fundamentally different from the instruction data of LLaMA-EΒ (Shi etΒ al., 2023) and EcomGPTΒ (Li etΒ al., 2023b), contains tasks all from real e-commerce platforms, and data all extracted from these real tasks.
@@ 3.2 Diverse Instructions
To test instruction understanding and model generalizability to new instructions, we hold out one instruction for each task as its ``unseen" instruction, which are inaccessible during model training. Thus, each task has 5 diverse instructions for instruction tuning, and 1 unseen instruction for testing. All the instructions are presented in AppendixΒ B.
@@ 3.4 Training, Validation, and Test Sets
is split into training sets, validation sets, IND test sets and OOD test sets as follows. TableΒ A2 summarizes the dataset. AppendixΒ A describes the details of the data split process.
@@
In-Domain Test SetΒ Β For each of the 10 tasks, also includes an in-domain (IND) test set of 1K samples (except for , which has 253 IND samples). We define IND in terms of products that are within the same category with the training products.
Out-of-Domain Test SetΒ Β To evaluate the generalizability of LLMs tuned on to unseen samples β a critical capacity desired to address the cold-start problem in e-commerce, we create OOD test sets in . We define OOD in terms of products, that is, new products unseen during LLM training are considered as OOD. We determine OOD products using their category information. In , 6 tasks β , , , , , and , have different product categories. For each of these tasks, all products from one particular category are held out as the OOD data for that task, as summarized in TableΒ A2 (each product belongs to only one category). We focus on OOD (new) products, because promoting and recommending new products to users are effective strategies in e-commcerce in driving sales and revenue, enhancing user experience, and increasing user engagement. Note we do not define OOD test sets over users (i.e., hold out users into OOD test sets), because users are anonymous and user identifiers are absent in existing datasets. Even though, the unseen instructions could be used in proximity to new users β a new user may have a completely new style of instructions to .
@@ 4 Models
We build a series of models on top of 6 base models: (1) : trained from large base models, Flan-T5 XXL (11B parameters)Β (Chung etΒ al., 2022) and Llama-2 13B-chatΒ (Touvron etΒ al., 2023b), (2) : trained from medium-sized base models, Llama-2 7B-chatΒ (Touvron etΒ al., 2023b) and Mistral-7B Instruct-v0.2Β (Jiang etΒ al., 2023), and (3) : trained from small base models, Flan-T5 XL (3B)Β (Chung etΒ al., 2022) and Phi-2 (3B)Β (Javaheripi & Bubeck, 2023). These models are instruction-tuned over training data.
We fine-tune all the base models with LoRAΒ (Hu etΒ al., 2021) and Huggingface transformers libraryΒ (Wolf etΒ al., 2019). During fine-tuning, the learning rate and batch size of all the models are set as 1e-4 and 128, respectively. We apply a cosine learning rate scheduler with a 5% warm-up period for 3 epochs. We set and the rank in LoRA as 16, and add LoRA adaptors to all the projection layers and the language modeling head. We perform 0-shot evaluation (i.e., without in-context examples) on all the tasks.

### v2

@@ Abstract
With tremendous efforts in developing effective e-commerce models, conventional e-commerce models show limited success in generalist e-commerce modeling, and suffer from unsatisfactory performance on new users and new products β a typical out-of-domain generalization challenge. Meanwhile, large language models (LLMs) demonstrate outstanding performance in generalist modeling and out-of-domain generalizability in many fields. Toward fully unleashing their power for e-commerce, in this paper, we construct , the first open-sourced, large-scale, and high-quality benchmark instruction dataset for e-commerce. Leveraging , we develop , a series of e-commerce LLMs, by instruction-tuning general-purpose LLMs. Our comprehensive experiments and evaluation demonstrate that models substantially outperform baseline models, including the most advanced GPT-4, and the state-of-the-art task-specific models in in-domain evaluation. Moreover, exhibits excellent generalizability to out-of-domain settings, including unseen products and unseen instructions, highlighting its superiority as a generalist e-commerce model. Both the dataset and the models show great potential in empowering versatile and effective LLMs for e-commerce. and models are publicly accessible through https://ninglab.github.io/eCeLLM/.
@@ 1 Introduction
The Internetβs evolution and the rise of the digital economy have made e-commerce an integral part of daily life, drawing considerable attention from researchers. In recent years, tremendous research efforts have been dedicated to developing effective e-commerce modelsΒ (Yang etΒ al., 2022; Geng etΒ al., 2022). Though promising, conventional e-commerce models generally suffer from two issues. (1) Limited success in generalist e-commerce modelingΒ (Zhang etΒ al., 2023b): Conventional e-commerce models are typically task-specific (e.g., gSASRec for sequential recommendationΒ (Petrov & Macdonald, 2023), SUOpenTag for attribute value extractionΒ (Xu etΒ al., 2019)). However, contemporary e-commerce platforms, such as Amazon and eBay, are highly complex and consistently expanding, with many interdependent tasks. In this context, generalist modeling is particularly suitable and desired on these platforms due to its cost-effectiveness and extensibility, while the task-specific modeling scheme struggles with scalability. (2) Unsatisfactory performance on new users and new productsΒ (Hou etΒ al., 2024): Cold start (i.e., performing e-commerce tasks on new users or new products)Β (Lika etΒ al., 2014), the typical out-of-domain (OOD) generalization challenge in e-commerce, has been a long-standing and difficult problemΒ (Ding etΒ al., 2022; Yang etΒ al., 2022). Existing task-specific e-commerce models are typically tailored to existing users and products and lack the ability to effectively extrapolate to new users and new productsΒ (Lika etΒ al., 2014). However, new users and new products are very common and highly wanted in the dynamic landscape of e-commerce.
Recently, large language models (LLMs), such as GPT-4Β (OpenAI, 2023), Claude 2Β (Anthropic, 2023), GeminiΒ (Team etΒ al., 2023), and Llama 2Β (Touvron etΒ al., 2023b), have demonstrated exceptional performance in natural language processingΒ (Zhao etΒ al., 2023), information retrievalΒ (Spatharioti etΒ al., 2023), and many other fieldsΒ (Frieder etΒ al., 2023; Zeng etΒ al., 2023). However, their power for e-commerce is not fully unleashed. While limited effortsΒ (Yang etΒ al., 2023; Zhang etΒ al., 2023a; Li etΒ al., 2024) have been dedicated to leveraging LLMs for e-commerce, most of them focus on studying the utility of pre-trained LLMs in single or homogenous tasks (e.g., e-commerce authoring). Some recent efforts employ instruction tuning to adapt LLMs for e-commerce applicationsΒ (Shi etΒ al., 2023; Li etΒ al., 2024; Geng etΒ al., 2022). However, they fall short in the covered e-commerce tasks and instruction data. We aim to bridge the gap and develop e-commerce foundation models with real-world utilities for a large variety of e-commerce applications.
@@
[…]
@@
FigureΒ 1 depicts the overall scheme of . To the best of our knowledge, this study is the first comprehensive and systematic study of instruction-tuning LLMs for e-commerce applications, open-sources the first large-scale, high-quality benchmark dataset (), and develops the state-of-the-art generalist LLMs ( series) for e-commerce. and models are publicly accessible through https://ninglab.github.io/eCeLLM/.
@@ 2 Related Work
Instruction Tuning with LLMsΒ Β Instruction tuning enables the transfer of general knowledge captured in LLMs to specific application domains, facilitating the generalizability of LLMs. FLANΒ (Wei etΒ al., 2021) and T0Β (Sanh etΒ al., 2021) are the early work to investigate instruction tuning by evaluating the zero-shot performance of fine-tuned LLMs on numerous datasets. Both of the models show encouraging performance over the original GPT-3Β (Brown etΒ al., 2020), indicating the effectiveness of instruction tuning. The importance of datasets, model scale, and instructions for instruction tuning are explored in a later work FLAN-v2Β (Chung etΒ al., 2024). Considering the significant impact of datasets for instruction tuning, several studies collect sizable benchmarks consisting of numerous tasks, such as Super-NaturalInstructionsΒ (Wang etΒ al., 2022b), Self-instructΒ (Wang etΒ al., 2022a) and Flan CollectionΒ (Longpre etΒ al., 2023), for fine-tuning general-purpose LLMs on NLP tasks. Meanwhile, instruction-tuned LLMs are employed in specific domains. For example, PlatypusΒ (Lee etΒ al., 2023a) is fine-tuned for reasoning in STEM, and MammothΒ (Yue etΒ al., 2024) is for general math problem-solving. Our incorporates instruction tuning with LLMs for e-commerce.
LLMs for E-commerceΒ Β LLMs are emerging in the general e-commerce realm. RecMindΒ (Wang etΒ al., 2023) utilizes LLMs as an autonomous recommender agent by enhancing their capability to comprehend user behaviors. LLaMA-EΒ (Shi etΒ al., 2023) is fine-tuned on LLaMAΒ (Touvron etΒ al., 2023a) for various e-commerce authoring tasks, such as ads generation and general product QA. EcomGPTΒ (Li etΒ al., 2024) is a pioneer instruction-tuned LLM on its dataset EcomInstruct. However, most of its tasks are human-constructed by repurposing data from other tasks (e.g., from the βquery product matchingβ task, a task is constructed to generate user queries based on the matched products), and thus, could have limited utility in real applications. A recent studyΒ (Geng etΒ al., 2022) develops a fine-tuned LLM, P5, with personalized prompts to conduct a few specific recommendation tasks (e.g., sequential recommendation). These studies demonstrate the utility of LLMs and instruction tuning in e-commerce applications. To the best of our knowledge, is the most comprehensive instruction-tuned e-commerce LLM.
Comparison among Instruction-tuned LLMs for E-commerceΒ Β TableΒ 1 summarizes the difference between our from LLaMA-E and EcomGPT, the two existing fine-tuned LLMs capable of performing multiple e-commerce tasks (P5 only performs recommendations). Fundamentally different from LLaMA-E and EcomGPT, focuses on comprehensive real-world e-commerce tasks and data, includes very extensive evaluation and benchmarking, and open-sources instruction data and models.
@@
| Β (Shi etΒ al., 2023) | Β (Li etΒ al., 2024) | (ours) | |
| Instruction tuning | β | β | β |
@@
| Real-world tasks | β | β | β |
| Β Β Β Β Β # training tasks | 5 | 122β‘ | 10 |
| Β Β Β Β Β # In-domain tests | 5 | 0 | 10 |
| Β Β Β Β Β # out-of-domain tests | 0 | 12β‘ | 6 |
| # general-purpose | 5 | 3 | 5 |
| LLMs evaluated | |||
| # task-specific SoTA models evaluated | 0 | 0 | 11 |
@@
β‘EcomGPT has 122 training tasks, most of which are manipulated from data of different other tasks. The training data is not publicly available. EcomGPT releases 12 test tasks (8 in Chinese) for only out-of-domain evaluation.
@@ 3 Dataset
We introduce , an instruction dataset for adapting LLMs to e-commerce tasks. features 3 key design principles. (1) Broad coverage: includes 10 diverse tasks of 4 categories. Comprehensive and wide-ranging related tasks are critical in enabling versatile LLMs for a specific domain as shown in the literatureΒ (Lee etΒ al., 2023b). (2) Realistic tasks: focuses on real-world e-commerce tasks with real-world data, not human-manipulated tasks with synthetic data. This ensures e-commerce LLMs tuned on a potentially high utility in real-world applications. (3) High quality: considers only real-world data and undergoes rigorous scrutiny to ensure its accuracy and high quality. High data quality plays a pivotal role in building effective LLMsΒ (Hoffmann etΒ al., 2022; Gadre etΒ al., 2024).
@@ 3.1 E-commerce Tasks
Particularly, includes 3 tasks for product understanding: (1) attribute value extraction ()Β (Xu etΒ al., 2019; Wang etΒ al., 2020; Yang etΒ al., 2022), (2) product matching ()Β (KΓΆpcke etΒ al., 2010; Rahm, Erhard, 2010) and (3) product relation prediction ()Β (Ahmed etΒ al., 2021; Xu etΒ al., 2020). For user understanding, includes (4) sentiment analysis ()Β (Wankhade etΒ al., 2022) and (5) sequential recommendation ()Β (Petrov & Macdonald, 2023; Li etΒ al., 2023). also covers 3 query product matching tasksΒ (Reddy etΒ al., 2022): (6) multi-class product classification (), (7) product substitute identification (), and (8) query-product ranking (). For product question answering, contains the tasks of (9) answerability prediction ()Β (Gupta etΒ al., 2019) and (10) answer generation ()Β (Deng etΒ al., 2023). More details are available in AppendixΒ A. TableΒ A1 summarizes the tasks and their data sources. Fundamentally different from the instruction data of LLaMA-EΒ (Shi etΒ al., 2023) and EcomGPTΒ (Li etΒ al., 2024), contains tasks all from real e-commerce platforms, and data all extracted from these real tasks.
@@ 3.2 Diverse Instructions
To test instruction understanding and model generalizability to new instructions, we hold out one instruction for each task as its βunseenβ instruction, which is inaccessible during model training. Thus, each task has 5 diverse instructions for instruction tuning, and 1 unseen instruction for testing. All the instructions are presented in AppendixΒ C.
@@ 3.4 Training, Validation, and Test Sets
is split into training sets, validation sets, IND test sets, and OOD test sets as follows. TableΒ A3 summarizes the dataset. AppendixΒ A describes the details of the data split process.
@@
In-Domain Test SetΒ Β For each of the 10 tasks, also includes an in-domain (IND) test set of 1K samples (except for , which has 253 IND samples). We define IND in terms of products that are within the same category as the training products.
Out-of-Domain Test SetΒ Β To evaluate the generalizability of LLMs tuned on to unseen samples β a critical capacity desired to address the cold-start problem in e-commerce, we create OOD test sets in . We define OOD in terms of products, that is, new products unseen during LLM training are considered as OOD. We determine OOD products using their category information. In , 6 tasks β , , , , , and , have different product categories. For each of these tasks, all products from one particular category are held out as the OOD data for that task, as summarized in TableΒ A3 (each product belongs to only one category). We focus on OOD (new) products, because promoting and recommending new products to users are effective strategies in e-commerce in driving sales and revenue, enhancing user experience, and increasing user engagement. Note we do not define OOD test sets over users (i.e., hold out users into OOD test sets), because users are anonymous and user identifiers are absent in existing datasets. Even though, the unseen instructions could be used in proximity to new users β a new user may have a completely new style of instructions to .
@@ 4 Models
We build a series of models on top of 6 base models: (1) : trained from large base models, Flan-T5 XXL (11B parameters)Β (Chung etΒ al., 2024) and Llama-2 13B-chatΒ (Touvron etΒ al., 2023b), (2) : trained from medium-sized base models, Llama-2 7B-chatΒ (Touvron etΒ al., 2023b) and Mistral-7B Instruct-v0.2Β (Jiang etΒ al., 2023), and (3) : trained from small base models, Flan-T5 XL (3B)Β (Chung etΒ al., 2024) and Phi-2 (3B)Β (Javaheripi & Bubeck, 2023). These models are instruction-tuned over training data.
We fine-tune all the base models with LoRAΒ (Hu etΒ al., 2021) and Huggingface transformers libraryΒ (Wolf etΒ al., 2019). During fine-tuning, the learning rate and batch size of all the models are set as 1e-4 and 128, respectively. We apply a cosine learning rate scheduler with a 5% warm-up period for 3 epochs. We set and the rank in LoRA as 16, and add LoRA adaptors to all the projection layers and the language modeling head. We perform 0-shot evaluations (i.e., without in-context examples) on all the tasks.